OpenStudy (anonymous):

How do I calculate the frequency in hertz of each wave. A: 419 NM 4.10X10^-7M B: 569NM 5.69X10^-7M C: 679NM 6.79X10^-7M PLEASE HELP ME

OpenStudy (anonymous):

These are all electromagnetic waves, right?

OpenStudy (anonymous):

\[ c = 3 x 10^8 m/s\] \[ c = f \lambda\] \[ f = \frac{c}{\lambda}\] Be sure to use the wavelength that's converted to meters. ^_^

OpenStudy (anonymous):

the Hertz unit is \[\frac{1}{s}\] a "per second"